OpenStudy (anonymous):

which ordered pair is a solution of this equation? 8x-5y=8 A. (0,2) B. (1,0) C. (2,0) D. (0,1)

OpenStudy (anonymous):

im not sure sorry

OpenStudy (anonymous):

its ok thanks for looking at it tho :)

OpenStudy (anonymous):

@MegänD

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

alright just give me a sec..:D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

its B. ...8x1-5x0=8

OpenStudy (anonymous):

thanks megan :)

OpenStudy (anonymous):

8x1=8-5x0=0 8-0=8 answer is 8

OpenStudy (anonymous):

your smart

OpenStudy (anonymous):

no bro im 16 so its easy for me...:D

OpenStudy (anonymous):

k

OpenStudy (anonymous):

:D have a good day you guys:D

OpenStudy (anonymous):

u too!

OpenStudy (anonymous):

u too:D